# Topological transition in cyanobacteria: from motion to structure

### v1

@@ III Conclusions and Outlook
We have studied the dynamics and structural properties of a colony of filamentous cyanobacteria as they undergo a topological transition to a reticulate pattern, a disordered network-like organization of filaments within bundles that link together. Filamentous cyanobacteria constitute a new class of active matter, which we have termed active spaghetti [70], due to the path-tracking nature of their gliding motility. We derived the nondimensional equations for a minimal model of their behavior that is firmly grounded in experimental evidence. The model captures the transient dynamics leading to the formation of the reticulate pattern. It also predicts a crossover in the mean-square displacement of filaments between ballistic and diffusive regimes, which can be rationalized in terms of an effective diffusivity. We explored the dynamics for different filament lengths, and found that the order parameter appears to approach a second order or continuous phase transition in the limit of very short filaments. The Péclet number dictates the typical length scale of the patterns, and together with the area coverage $\Phi$ describes the relevant fields of the nonequilibrium phase diagram for our system. Inspired by polymer physics, we introduced parallel and perpendicular pair correlation functions, which capture the fundamental structural properties of the reticulate bundles. In particular, the minimum of $g_{\perp}(r)$ faithfully predicts the typical length scale of the pattern, $\ell^{*}$. These observations can help further understand the formation and properties of cyanobacterial biomats.
@@
Figure 10: O. lutea grown on agar for six months shows a more complex, three-dimensional structure. Photographs of the top surface of the biomat (left panel: face-on view; right panel: tilted view) shows a complex topography of raised, filled circular features, which develop out of an earlier template of a more open, reticulate structure.
@@
Biofilms and biomats are self-organized microbial communities that differentiate and develop phenotypic adaptation both in space and time [8]. It stands to reason that the form of the reticulate patterns studied here has a biological function sculpted by survival needs. The reticulate network then serves as a foundation for mature biofilms that adapts to the features of the surrounding environment. Because gliding cyanobacteria are surface associated, a natural question is to inquire how the surface properties influence the growth and maturation of the biofilm [88, 89]. As an initial exploration of the effects of substrate stiffness on colony formation, we also cultivated O. lutea on agar plates, see Fig. 10. These were prepared by mixing 1.5% agar, by weight, with a 1:50 dilution of BG11 broth in deionized water, following the protocols of Ref. [90]. Every eight weeks, 2 ml of 1:100 BG11 broth were added to the plates. Samples were otherwise left undisturbed, growing under identical conditions to those described for the transient experiment (Fig. 2), and following Ref. [70]. Over time, the initially reticulate pattern developed into a three-dimensional structure, with the open areas of the network filling in and growing upward that was templated on the original network boundaries. These observations point at the physical richness of this biological system and invite future exploration.
IV Acknowledgments
@@
We thank Maike Lorenz (SAG Göttingen) for support with cyanobacteria cultures, Stefan Karpitschka from Max Planck Institute for Dynamics and Self-organisation (MPIDS) and Jack Paget (Loughborough) for discussions, and Graham Hickman (NTU) for microscopy support. We gratefully acknowledge Ellen Nisbet (University of Nottingham) for helpful discussions. Microscopy facilities were provided by the Imaging Suite at the School of Science and Technology at Nottingham Trent University. Numerical calculations were performed using the Sulis Tier 2 HPC Platform funded by EPSRC Grant EP/T022108/1 and the HPC Midlands+ consortium. We gratefully acknowledge use of the Lovelace HPC service at Loughborough University. M.K.F. was partly sponsored by the Malawi University of Science and Technology. This work was supported by the MPIDS. J.C. was supported by EPSRC grant EP/W522569/1 and UKRI/Wellcome grant no EP/T022000/1–PoLNET3. N.D. was supported by a BBSRC doctoral training programme (BB/T0083690/1).
V Author contributions
@@ VI Competing interests
VII Code availability
The simulation code is available upon reasonable request.

### v2

@@ IV Conclusions
Filamentous cyanobacteria, which we have termed active spaghetti [72], constitute a new class of active matter, due to the path-tracking nature of their gliding motility. We derived the nondimensional equations for a minimal model of their behavior that is firmly grounded in experimental evidence. The model captures the transient dynamics leading to the formation of the reticulate pattern. It also predicts a crossover in the mean-square displacement of filaments between ballistic and diffusive regimes, which can be rationalized in terms of an effective diffusivity. We explored the dynamics for different filament lengths, and found that the order parameter appears to approach a second order or continuous phase transition in the limit of very short filaments. This is in contrast to the well-studied Vicsek model [16] which may exhibit a weak-first order transition [99]. Furthermore, our system does not exhibit an ordered phase where all filaments point in the same direction, on account of the curvature fluctuations present in our model. The Péclet number dictates the typical length scale of the patterns, and together with the area coverage $\Phi$ describes the relevant fields of the nonequilibrium phase diagram for our system. Inspired by polymer physics, we introduced parallel and perpendicular pair correlation functions, which capture the fundamental structural properties of the reticulate bundles. In particular, the minimum of $g_{\perp}(r)$ faithfully predicts the typical length scale $\ell^{*}$ of the pattern. These observations can help further understand the formation and properties of cyanobacterial biomats.
@@
Figure 11: Mature cyanobacteria biofilm. O. lutea grown on agar for six months shows a more complex, three-dimensional structure. Photographs of the top surface of the biomat (left panel: face-on view; right panel: low-angle shot of the same sample) shows a complex topography of raised, filled circular features, which develop out of an earlier template of a more open, reticulate structure.
@@
Biofilms and biomats are self-organized microbial communities that differentiate and develop phenotypic adaptation both in space and time [8]. It stands to reason that the form of the reticulate patterns studied here has a biological function sculpted by survival needs. The reticulate network then serves as a foundation for mature biofilms that adapts to the features of the surrounding environment. Even growing as films on glass culture bottles, these biofilms can easily reach 100 $\mu$m thick, much thicker than an individual filament. Because gliding cyanobacteria are surface associated, a natural question is then to inquire how the surface properties influence the growth and maturation of the biofilm [100, 101]. As an initial exploration of the effects of substrate stiffness on colony formation, we also cultivated O. lutea on agar plates, see Fig. 11. These were prepared by mixing 1.5% agar, by weight, with a 1:50 dilution of BG11 broth in deionized water, following the protocols of Ref. [102]. Every eight weeks, 2 ml of 1:100 BG11 broth were added to the plates. Samples were otherwise left undisturbed, growing under identical conditions to those described for the transient experiment (Fig. 2), and following Ref. [72]. Over time, the initially reticulate pattern developed into a three-dimensional structure, with the open areas of the network filling in and growing upward that was templated on the original network boundaries. These observations point at the physical richness of this biological system and invite future exploration.
V Acknowledgments
@@
We thank Maike Lorenz (SAG Göttingen) for support with cyanobacteria cultures, Stefan Karpitschka from the Max Planck Institute for Dynamics and Self-organisation (MPIDS) and Jack Paget (Loughborough University) for discussions, and Graham Hickman at Nottingham Trent University (NTU) for microscopy support. We gratefully acknowledge Ellen Nisbet (University of Nottingham) for helpful discussions, and Alex Luke (NTU) for preliminary work exploring the imaging of biofilms. Microscopy facilities were provided by the Imaging Suite at the School of Science and Technology at NTU. Numerical calculations were performed using the Sulis Tier 2 HPC Platform funded by EPSRC Grant EP/T022108/1 and the HPC Midlands+ consortium. We gratefully acknowledge use of the Lovelace HPC service at Loughborough University. M.K.F. was partly sponsored by the Malawi University of Science and Technology. This work was supported by the MPIDS. J.C. was supported by EPSRC grant EP/W522569/1 and UKRI/Wellcome grant EP/T022000/1–PoLNET3. N.D. was supported by a BBSRC doctoral training programme (BB/T0083690/1).
VI Author contributions
@@ VII Competing interests
VIII Data availability
All data needed to evaluate the conclusions in the paper are present in the paper. All micrographs used are shown in the paper. Micrographs in their original resolution are available upon reasonable request to L.G.
IX Code availability
The simulation code and output data are available upon reasonable request to M.G.M.